OpenStudy (anonymous):

What is the vertex for the function? (You will want to find the axis of symmetry first) Write as an ordered pair leaving no spaces, such as, (1,-4) y=-3x^2+6x

OpenStudy (amistre64):

given: ax^2+bx+c the vertex is on the axis of symmetry: -b/2a when x= -b/2a we find y

OpenStudy (amistre64):

in this case -3x^2 +6x +0 gives us: a=-3 b=6

OpenStudy (amistre64):

oooo, -1 :)

OpenStudy (amistre64):

err, +1

OpenStudy (amistre64):

-6/2(-3) = -6/-6 = 1

OpenStudy (amistre64):

so when x=1; we get y= -3+6